# FonTS: Text Rendering with Typography and Style Controls

### v2

@@ Abstract
Visual text rendering are widespread in various real-world applications, requiring careful font selection and typographic choices. Recent progress in diffusion transformer (DiT)-based text-to-image (T2I) models show promise in automating these processes. However, these methods still encounter challenges like inconsistent fonts, style variation, and limited fine-grained control, particularly at the word-level. This paper proposes a two-stage DiT-based pipeline to address these problems by enhancing controllability over typography and style in text rendering. We introduce typography control fine-tuning (TC-FT), an parameter-efficient fine-tuning method (on $5\%$ key parameters) with enclosing typography control tokens (ETC-tokens), which enables precise word-level application of typographic features. To further address style inconsistency in text rendering, we propose a text-agnostic style control adapter (SCA) that prevents content leakage while enhancing style consistency. To implement TC-FT and SCA effectively, we incorporated HTML-render into the data synthesis pipeline and proposed the first word-level controllable dataset. Through comprehensive experiments, we demonstrate the effectiveness of our approach in achieving superior word-level typographic control, font consistency, and style consistency in text rendering tasks. The datasets and models will be available for academic use.
@@ 1 Introduction
Visual text images are ubiquitous in daily life and hold significant commercial value in advertising, branding, and marketing [4, 10]. However, the design process for visual text is complex and time-consuming. Designers must carefully select appropriate fonts, use typographic elements like italics, and create artistic styles that are aesthetically pleasing and coherent. Recent advances in diffusion models [35, 31] demonstrate promising potential for creating visual contents in design, thereby attracting substantial attention. Concurrently, real-world applications raise increasing demands for controllability over the generated content.
Previous efforts have mainly focused on improving control over the content accuracy of scene text rendering [48, 42, 9, 10]. With the development of DiT-based T2I models, e.g. SD3 [15] and Flux .1 [1], the accuracy of text content has seen significant improvements. Beyond content accuracy, Glyph-ByT5 [23] introduced a new text encoder through contrastive learning, enabling various font types of text. Textdiffuser-2 [10] trained both two language models and the whole diffusion model to acquire layout planing capabilities. While these methods [23, 10] have implemented control at the paragraph-level, no methods have yet realized word-level control. Moreover, prior methods often overlook the artistic aspects of text  [10]. Recent DiT models [15, 1] have demonstrated promising capabilities in artistic text rendering, yet they face challenges like semantic confusion and style inconsistency.
@@
To this end, we propose a two-stage DiT-based pipeline for text rendering with typography and style controls. For typography control, we introduce Typography Control (TC)-finetuning, an parameter-efficient fine-tuning method, alongside enclosing typography control tokens (ETC-tokens). By introducing HTML-render to ingeniously design the data synthesis pipeline, we propose the first word-level typography control dataset (TC-dataset). Our findings show that the model not only learns typographic elements but also applies specific typographic features at precise word locations. For style control, we introduces a style control adapter (SCA) that injects style information without compromising the accuracy of the text. The training of SCA is also a two-stage process, each stage using a different dataset. In total, these datasets consist of approximately 600k image-text pairs with high aesthetic scores.
@@
  • We introduce the first word-level controllable dataset. By leveraging ETC-tokens, we enables precise learning of typographic attributes and their specific locations.
@@ 2 Related Work
Scene Text Rendering. Despite progress in diffusion models [35, 31], high-quality scene text rendering remains a challenge. To address this, one line of research [48, 10, 9, 41] focuses on explicitly controlling the position and content of the text being rendered, relying on ControlNet [51]. Another line of works [23, 24] fine-tune the character-aware ByT5 text encoder [22] using paired glyph-text datasets, improving the ability to render accurate text in images.
Artistic Text Rendering. Early research focused on font creation by transferring textures from existing characters, employing stroke-based methods [6], patch-based techniques [44, 45, 46], and GAN-based [3, 18, 17, 47, 25] methods. Innovations with diffusion models [38, 43, 28] have enabled diverse text image stylization and semantic typography, resulting in visually appealing designs that retain readability. However, despite recent DiT models [15, 1] showing quite promise in artistic text rendering, they still struggle with semantic confusion and style inconsistency.
Controllable Image Generation. Text-based controllable methods [16, 36, 19] customize image outputs by fine-tuning diffusion models using user-provided examples. These approaches introduce modifier tokens to guide the generation process. ColorPeel [7] further enhances this by constructing datasets of color-shape pairs, to generate images with target colors. Image-based controllable methods, such as UniControl [32] retrain T2I models from scratch, are computationally expensive [51]. A efficient alternative introduces trainable modules to existing architectures [27, 49, 51]. These adapters enable structural  [51, 27] and style control from images [49, 8].
Most prior approaches are implemented on U-Net with a single CLIP text encoder. There has been relatively limited exploration of DiT-based T2I models that can incorporate multiple text encoders. Moreover, the area of controllable generation under multi-modal conditions has not been well-explored. Our work extends them to DiT-based models, enabling word-level typographic control, a more fine-grained form of control than previously achieved. Additionally, we seamlessly integrate both text-based and image-based controls to expand the range of real-world applications.
@@ 4.1 Quantitative Results
User Studies. We conducted user studies to perceptually evaluate our results against baselines. The evaluation centered on two key aspects: font consistency (Font-Con) and style consistency (Style-Con). Details are provided in the supplementary Sec 7 .
@@ 4.3 Ablation study
Ablation on SCA. Our style control adapters (SCA) are trained through two phases as mentioned in Sec  3.2, and ablation is focus on the second phase. Comparing top two rows in Figure 7, it is evident that TC-FT enhances text accuracy, yet severely weakens the artistry. Shifting to the third row, Art-FT significantly boosts artistry without damaging the accuracy. The fourth row, being nearly identical to the third, suggests that after Art-FT, TC-FT have a minimal negative impact on artistry. Results presented in Table 7 further validate this observation within the ATR task, and we also compared SCA on MM-DiT only with SCA on MM-DiT and Single-DiT both. The results are detailed in supplementary Sec 2.1.
Ablation on ETC-Tokens. The ETC tokens are designed for assigning the words which need to be controlled. We consider three cases: 1) Non-token: directly use the prompt as “make the ‘word’ Bold”; 2) single token: use single token in front of the ‘word’ same in  [19, 7]; 3) Ours. The results are detailed in supplementary Sec 2.3.
@@ 4.4 Applications and Limitation
Applications. Artistic font design. Benefit from the robust style consistency, our approach is able to generate a variety of artistic letters with high consistency. Moreover, because the SCA is pre-trained on high-quality, large-scale data, the style control is not limited to artistic text images. Any style image can be used as a control input, as shown in Figure 8.
Logo design. Scene text and artistic text images can also be seamlessly integrated. By using scene text image prompts alongside artistic text images for style control, our method achieves a smooth blend of the two, as shown in Figure 9. This allows for the creation of versatile logo designs that are suitable for a range of application scenarios.
@@ 5 Conclusion
We propose a two-stage DiT-based pipeline for text rendering with typography and style controls. TC-FT with ETC-tokens enables the model to learn and apply word-level attributes. The style control adapter facilitates style control without compromising text content. Additionally, we introduce the first word-level control dataset. Experimental results demonstrate that our method outperforms baselines in font consistency and style consistency, and word-level controls for text rendering tasks. This paper is the first to achieve word-level control in text rendering, in future work, we plan to explore its extension to multilingual rendering.
@@
Figure 22: Ablation study of style control adapter (SCA), results from style captions only after 10k and 40k steps of TC-finetunig.
@@
Figure 25: Results of fine-tuing T5 text encoder with new tokens, while input for CLIP is fix prompt: ‘words only, clean background’.
Figure 26: Results of our method: (a), (b) and (c) in basic text rendering, artistic text rendering, and scene text rendering, respectively.
@@ A.2 Differences with Flux-IPA
1) Our style control adapters (SCA) employ a two-stage training approach. Fine-tuning with SC-artext significantly boosts artistry without compromising the accuracy of text, making it more suitable for ATR task.
2) In contrast to Flux-IPA(XLabs)  [^5] Flux-IPA(XLabs), which is only applied on MM-DiT, our SCA is implemented on both MM-DiT and Single-DiT to enhance style control, as depicted in Figure 13 with Figure  14. Even with a style image scale of 0.6, the style achieved by applying SCA on both MM-DiT and Single-DiT is markedly superior to that of applying SCA only on MM-DiT with a style image scale of 0.9. The comparison between Table 7 and Table 8 further validates this, as applying SCA on both on MM-DiT and Single-DiT yields a higher CLIP-I score under different settings.
3) Unlike Flux-IPA(InstantX)  [^6] Flux-IPA(InstantX) which uses SigLIP [50], our method select CLIP [33] as the image encoder. This choice is grounded in the distinct characteristics of these two models. SigLIP [50, 40] is renowned for its robust OCR capabilities. Conversely, as discussed in [23, 12], CLIP’s visual embeddings are insensitive to text. This insensitivity to text in CLIP’s visual embeddings is pivotal for our application, as it mitigates content leakage from style images (artistic text images). The visual outcomes presented in Figure 15 provide empirical evidence support for our selection.
4) Distinct from previous methods, we insert adapters in an interval-skip manner (on layer 0,2,4…) to reduce costs. In terms of parameter usage, the parameters of adapters in Flux-IPA(InstantX) is approximately 2.85 times of ours, as demonstrated in Table 9.
@@ A.3 More Qualitative Results of ATR
This section serves as a supplement to Section 4.2 of the main paper, offering a qualitative comparison of our method with Glyph-ByT5 [23] and Textdiffuser-2 [10] on the ATR-bench dataset, as depicted in Figure 16. Additionally, we present our extended qualitative results on the ATR-bench dataset, including single-word and multi-word examples, are presented in Figure 17. Notably, in the second row of results in Figure 17, the accurate mirror reflection of letters in the every results further substantiates the effectiveness of our SCA. This example showcases that our SCA can inject style while meticulously maintaining text accuracy, providing additional empirical evidence for the capabilities of our proposed approach in the ATR task.
@@ A.4 Train Baseline
In addition to the aforementioned comparisons, we fine-tune another baseline, AnyText [41] on the TC-dataset using a method similar to TC-finetuning. The quantitative results are presented in Table 11, while the qualitative results are shown in Figure 18. These results clearly reveals that AnyText fails to acquire word-level controllability. The performance of Glyph-ByT5 and Textdiffuser-2 exhibits similar limitations. This may be attributed to the inherent restricted capabilities of the base models for text rendering. Figure 19 shows the attention maps of different base models for different words in basic text rendering.
@@ B.1 Ablation on SCA
SCA with Art-FT and TC-FT. The CLIP-I and OCR-Acc presented in Table 7 are the average figures obtained on ATR task when the scale of the style image is set at 0.9 and 0.6 respectively. Table 7 is identical to Table 6 in the main paper. These values are placed here to enable a more direct comparison with SCA only on MM-DiT (Table 8). It becomes evident that, irrespective of whether SCA, the impacts of Art-FT and TC-FT on the ATR task remain consistent: Art-FT enhances stylization, while TC-FT improves content accuracy. Additionally, as shown in Table 10, after Art-FT, the degree of style degradation caused by TC-FT is reduced. This highlights the distinct but complementary roles of Art-FT and TC-FT in optimizing both the stylistic and content-related aspects of the results.
@@ B.3 Ablation on ETC-Tokens
This section supplements Section 4.3 of the main paper, focusing on demonstrating the effectiveness of the proposed Enclosing Typography Control (ETC)-tokens for targeted word-level typographic attributes. For instance, to bold the word ”robot” in the phrase “i am not a robot”, we explore three settings: 1) Non-Token: Using a instruction prompt instead of adding modifier tokens, such as “Make ‘robot’ bold”. 2) Single-Token: Following [19, 7], we trained our model to use a single token, placing the modifier token before “robot”. 3) Our ETC-Token. The visual results of ablation on ETC-tokens are presented in Figure 21.
@@ Appendix E Details of Datasets
SC-artext. For fine-tuning the style control adapters, we created the SC-artext dataset. We combined a list of 100 style descriptions with a list of 99 words, categorized into three character length groups: 1-15, 16-30, and 30-50. This combination produced a variety of prompts for artistic text images, which served as input for Flux.1-dev [1], yielding around 20k high-quality images. To ensure the images accurately reflected the original text content, we utilized shareGPT4v [11] to regenerate captions. Figure 28 (b) shows sample images, and Table 13 presents the paired texts.

### v3

@@ Abstract
Visual text rendering is widespread in various real-world applications, requiring careful font selection and typographic choices. Recent progress in diffusion transformer (DiT)-based text-to-image (T2I) models shows promise in automating these processes. However, these methods still encounter challenges like inconsistent fonts, style variation, and limited fine-grained control, particularly at the word-level. This paper proposes a two-stage DiT-based pipeline to address these problems by enhancing controllability over typography and style in text rendering. We introduce typography control fine-tuning (TC-FT), a parameter-efficient fine-tuning method (on $5\%$ key parameters) with enclosing typography control tokens (ETC-tokens), which enables precise word-level application of typographic features. To further address style inconsistency in text rendering, we propose a text-agnostic style control adapter (SCA) that prevents content leakage while enhancing style consistency. To implement TC-FT effectively, we incorporated an HTML-rendered data pipeline and proposed the first word-level controllable dataset. Through comprehensive experiments, we demonstrate the effectiveness of our approach in achieving superior word-level typographic control, font consistency, and style consistency in text rendering tasks. Our project page is available at this site.
@@ 1 Introduction
Visual text images are ubiquitous in daily life and hold significant commercial value in advertising, branding, and marketing [4, 10]. However, the design process for visual text is complex and time-consuming. Designers must carefully select appropriate fonts, use typographic elements like italics, and create artistic styles that are aesthetically pleasing and coherent. Recent advances in diffusion models [51, 47] demonstrate promising potential for creating visual content in design, thereby attracting substantial attention. Concurrently, real-world applications raise increasing demands for control over the generated content [28, 54, 87].
Previous efforts have mainly focused on improving control over the content accuracy of scene text rendering [78, 68, 9, 10]. With the development of DiT-based T2I models, e.g., SD3 [16] and Flux.1 [1], the accuracy of text content has seen significant improvements. Beyond content accuracy, Glyph-ByT5 [33] introduced a new text encoder through contrastive learning [88], enabling various font types of text. Textdiffuser-2 [10] trained both two language models and the whole diffusion model to acquire layout planning capabilities. While these methods [33, 10] have implemented control at the paragraph-level, no methods have yet realized word-level control. Moreover, prior methods often overlook the artistic aspects of text  [10]. Recent DiT models [16, 1] have demonstrated promising capabilities in artistic text rendering, yet they face challenges like semantic confusion and style inconsistency.
@@
To this end, we propose a two-stage DiT-based pipeline for text rendering with typography and style controls. For typography control, we introduce Typography Control (TC)-finetuning, a parameter-efficient fine-tuning method, alongside enclosing typography control tokens (ETC-tokens). By introducing HTML-render to ingeniously design the data synthesis pipeline, we propose the first word-level typography control dataset (TC-dataset). Our findings show that the model not only learns typographic elements but also applies specific typographic features at precise word locations. For style control, we introduce a style control adapter (SCA) that injects style information without compromising the accuracy of the text. The training of SCA is also a two-stage process, each stage using a different dataset. In total, these datasets consist of approximately 600k image-text pairs with high aesthetic scores.
@@
  • We introduce the first word-level controllable dataset. By leveraging ETC-tokens, we enable precise learning of typographic attributes and their specific locations.
@@ 2 Related Work
Scene Text Rendering. Despite progress in diffusion models [51, 47], high-quality scene text rendering remains a challenge. To address this, prior research [78, 10, 9, 67] focuses on explicitly controlling the position and content of the text being rendered, relying on ControlNet [81]. Another line of works [33, 34] fine-tunes the character-aware ByT5 text encoder [32] using paired glyph-text datasets, improving the ability to render accurate text in images.
Artistic Text Rendering. Early research focused on font creation by transferring textures from existing characters, employing stroke-based methods [6, 55, 56], patch-based techniques [74, 75, 76], and GAN-based [3, 25, 21, 77, 41, 63, 73] methods. Innovations with diffusion models [64, 70, 44, 53, 35] have enabled diverse text image stylization and semantic typography, resulting in visually appealing designs that retain readability. However, despite recent DiT models [16, 1] showing quite promise in artistic text rendering, they struggle with semantic confusion and style inconsistency.
Controllable Image Generation. Controllable generation methods [57, 82, 85, 83, 71, 13, 37, 39, 36, 38, 40, 90, 89] enable control over diffusion models to synthesize specific subjects or layouts, typically by fine-tuning on user-provided examples and modifying the attention mechanism. For text-based controllable methods [20, 52, 26], ColorPeel [7] constructs color-shape pairs to generate images with target colors. For image-based controllable methods, UniControl [48] retrain T2I models from scratch, which is computationally expensive [81]. A more efficient alternative involves integrating trainable modules into existing models as adapters, enabling structural  [79, 43, 30] and style controls [8, 22, 86, 29].
Most prior approaches are implemented on U-Net with a single CLIP text encoder. While existing methods have explored controllable generation under the Diffusion Transformer (DiT) architecture  [84, 62, 58, 59, 24, 60, 23, 72, 69], such studies remain relatively limited. Moreover, the area of controllable generation under multi-modal conditions has not been well-explored. Our work advances this direction by enabling word-level controls and seamlessly integrating multi-modal controls to broaden applications.
@@ 4.1 Quantitative Results
User Studies. We conducted user studies to perceptually evaluate our results against baselines on two key aspects: font consistency (Font-Con) and style consistency (Style-Con). Details are provided in the supplementary Sec 7 .
@@ 4.3 Ablation study
Ablation on SCA. Style control adapters (SCA) are trained through two phases as mentioned in Sec 3.2, and ablation focuses on the second phase. Comparing top two rows in Figure 7, it is evident that TC-FT enhances text accuracy, yet severely weakens the artistry. Shifting to the third row, Art-FT significantly boosts artistry without damaging the accuracy. The fourth row, being nearly identical to the third, suggests that after Art-FT, TC-FT has a minimal negative impact on artistry. Results presented in Table 7 further validate this observation in ATR, and we also compared SCA on MM-DiT only, with MM-DiT and Single-DiT both. The results are detailed in supplementary Sec 2.1.
Ablation on ETC-Tokens. The ETC tokens are designed for assigning the words which need to be controlled. We consider three cases: 1) Non-token: directly use the prompt as “the ‘word’ in Bold”; 2) single token: use a single token; 3) Ours. The results are detailed in supplementary Sec 2.3.
@@ 4.4 Applications and Limitation
Applications. Artistic font design. Benefit from the robust style consistency, our approach can generate a variety of artistic letters with high consistency. Moreover, because the SCA is pre-trained on high-quality, large-scale data, the style control is not limited to artistic text images. Any style image can be used as a control input, as shown in Figure 8.
Logo design. Scene text and artistic text images can also be seamlessly integrated. By using scene text image prompts alongside artistic text images for style control, our method achieves a smooth blend of the two, as shown in Figure 9. These show our method is suitable for various applications.
@@ 5 Conclusion
We propose a two-stage DiT-based pipeline for text rendering with typography and style controls. TC-FT with ETC-tokens enables the model to learn and apply word-level attributes. The style control adapter facilitates style control without compromising text content. Additionally, we introduce the first word-level control dataset. Experimental results demonstrate that our method outperforms baselines in font consistency and style consistency, and word-level controls for text rendering tasks. This paper is the first to achieve word-level control in text rendering. In future work, we plan to explore its extension to multilingual rendering.
Acknowledgment
This work was substantially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/RGC Project PolyU 25211424) and partially supported by a grant from PolyU university start-up fund (Project No. P0047675).
@@ Acknowledgment
Figure 22: Infer without ‘sks’ to mitigate scene-text detachment.
Figure 23: Ablation study of style control adapter (SCA), results from style captions only after 10k and 40k steps of TC-finetuning.
@@
Figure 26: Results of fine-tuning T5 text encoder with new tokens, while input for CLIP is fixed prompt: ‘words only, clean background’.
Figure 27: Results of our method: (a), (b), and (c) in basic text rendering, artistic text rendering, and scene text rendering, respectively.
@@
Figure 30: Visual comparison of existing artistic text datasets.
@@ A.2 Differences with Flux-IPA
1) Our style control adapters (SCA) employ a two-stage training approach. Fine-tuning with SC-artext significantly boosts artistry without compromising the accuracy of text, making it more suitable for the ATR task.
2) In contrast to Flux-IPA(XLabs)  [^5] Flux-IPA(XLabs), which is only applied on MM-DiT, our SCA is implemented on both MM-DiT and Single-DiT to enhance style control, as depicted in Figure 13 with Figure 14. Even with a style image scale of 0.6, the style achieved by applying SCA on both MM-DiT and Single-DiT is markedly superior to that of applying SCA only on MM-DiT with a style image scale of 0.9. The comparison between Table 7 and Table 8 further validates this, as applying SCA on both MM-DiT and Single-DiT yields a higher CLIP-I score under different settings.
3) Unlike Flux-IPA(InstantX)  [^6] Flux-IPA(InstantX) which uses SigLIP [80], our method select CLIP [49] as the image encoder. This choice is grounded in the distinct characteristics of these two models. SigLIP [80, 66] is renowned for its robust OCR capabilities. Conversely, as discussed in [33, 12], CLIP’s visual embeddings are insensitive to text. This insensitivity to text in CLIP’s visual embeddings is pivotal for our application, as it mitigates content leakage from style images (artistic text images). The visual outcomes presented in Figure 15 provide empirical evidence in support of our selection.
4) Distinct from previous methods, we insert adapters in an interval-skip manner (on layer 0,2,4…) to reduce costs. In terms of parameter usage, the parameters of adapters in Flux-IPA(InstantX) are approximately 2.85 times ours, as demonstrated in Table 9.
@@ A.3 More Qualitative Results of ATR
This section serves as a supplement to Section 4.2 of the main paper, offering a qualitative comparison of our method with Glyph-ByT5 [33] and Textdiffuser-2 [10] on the ATR-bench dataset, as depicted in Figure 16. Additionally, we present our extended qualitative results on the ATR-bench dataset, including single-word and multi-word examples, in Figure 17. Notably, in the second row of results in Figure 17, the accurate mirror reflection of letters in every result further substantiates the effectiveness of our SCA. This example showcases that our SCA can inject style while meticulously maintaining text accuracy, providing additional empirical evidence for the capabilities of our proposed approach in the ATR task.
@@ A.4 Train Baseline
In addition to the aforementioned comparisons, we fine-tune another baseline, AnyText [67] on the TC-dataset using a method similar to TC-finetuning. The quantitative results are presented in Table 11, while the qualitative results are shown in Figure 18. These results clearly reveal that AnyText fails to acquire word-level controllability. The performance of Glyph-ByT5 and Textdiffuser-2 exhibits similar limitations. This may be attributed to the inherent restricted capabilities of the base models for text rendering. Figure 19 shows the attention maps of different base models for different words in basic text rendering.
@@ B.1 Ablation on SCA
SCA with Art-FT and TC-FT. The CLIP-I and OCR-Acc presented in Table 7 are the average figures obtained on ATR task when the scale of the style image is set at 0.9 and 0.6, respectively. Table 7 is identical to Table 6 in the main paper. These values are placed here to enable a more direct comparison with SCA only on MM-DiT (Table 8). It becomes evident that, irrespective of whether SCA, the impacts of Art-FT and TC-FT on the ATR task remain consistent: Art-FT enhances stylization, while TC-FT improves content accuracy. Additionally, as shown in Table 10, after Art-FT, the degree of style degradation caused by TC-FT is reduced. This highlights the distinct but complementary roles of Art-FT and TC-FT in optimizing both the stylistic and content-related aspects of the results.
@@ B.3 Ablation on ETC-Tokens
This section supplements Section 4.3 of the main paper, focusing on demonstrating the effectiveness of the proposed Enclosing Typography Control (ETC)-tokens for targeted word-level typographic attributes. For instance, to bold the word ”robot” in the phrase “i am not a robot”, we explore three settings: 1) Non-Token: Using an instruction prompt instead of adding modifier tokens, such as “the ‘robot’ is in bold”. 2) Single-Token: Following [26, 7], we trained our model to use a single token, placing the modifier token before “robot”. 3) Our ETC-Token. The visual results of ablation on ETC-tokens are presented in Figure 21.
B.4 Ablation on ‘sks’ prefix
We tried to mitigate language drift (scene-text detachment) by training with the ‘sks’ prefix in prompts of the TC-Dataset, which is omitted during inference. This low-cost approach helps alleviate detachment, as shown in Figure 22.
@@ Appendix E Details of Datasets
SC-artext. For fine-tuning the style control adapters, we created the SC-artext dataset. We combined a list of 100 style descriptions with a list of 99 words, categorized into three character length groups: 1-15, 16-30, and 30-50. This combination produced a variety of prompts for artistic text images, which served as input for Flux.1-dev [1], yielding around 20k high-quality images. To ensure the images accurately reflected the original text content, we utilized shareGPT4v [11] to regenerate captions. Figure 29 (b) shows sample images, and Table 14 presents the paired texts. Besides, we provide quantitative and qualitative comparison results with artistic text in TWD [18] and Posta [17] in Table 13 and Figure 30, respectively. We randomly sample 100 images from each and use specialized LMM (Q-Align [19]) for quality and aesthetic evaluation.